OpenStudy (anonymous):

differential equation SOS y'-y=2x-3

OpenStudy (freckles):

this is a linear differentiation equation that can be solved with a integrating factor

OpenStudy (anonymous):

I don't get to the final answer

OpenStudy (anonymous):

yeah, it's just the beginning

OpenStudy (freckles):

\[v=e^{\int\limits (-1) dx}\] this is the integrating factor -1 is up there becaue your diff is in the form y'+py=q where we choose the integrating factor to be \[v=e^{\int\limits p dx}\]

OpenStudy (freckles):

so can you integrate the -1 up there w.r.t. x?

OpenStudy (anonymous):

we didn't learn that....

OpenStudy (freckles):

oh then what are you guys doing

OpenStudy (freckles):

like are you doing that like guessing type thing where you try to determine some coefficients bases on what you think the solution will look like or whatever they call it

OpenStudy (anonymous):

Singular solutions

OpenStudy (anonymous):

like putting dy/dx instead of y'

OpenStudy (freckles):

I don't think singular solutions is a method to solve the differential equation

OpenStudy (anonymous):

maybe first order?

OpenStudy (freckles):

I don't get it. Are we not suppose to solve the differential equation?

OpenStudy (anonymous):

idk the title tbh, our teacher doesn't speak much the language, he does solve though

OpenStudy (anonymous):

is there may be an easier way?

OpenStudy (freckles):

so if he said to solve then you must know a method to solve this first order linear differential equation

OpenStudy (anonymous):

It was our first lesson after all

OpenStudy (freckles):

and I need to know the method

OpenStudy (anonymous):

well thanks though

OpenStudy (freckles):

I know two methods. One is called using a integrating factor and the other is called something about undetermined coefficients. So neither of those ring a bell for you?

OpenStudy (anonymous):

the first one yes!

OpenStudy (anonymous):

out TA told us to switch y' with dy/dx

OpenStudy (freckles):

oh i thought you said you didn't know that method earlier

OpenStudy (freckles):

I was doing the whole integrating factor thing and you know you didn't know that route remember?

OpenStudy (anonymous):

I'll go though on it again @freckles by myself thanks though

OpenStudy (anonymous):

I'll try again by myself first I think

OpenStudy (freckles):

\[y'+py=q \\ \text{ Integrating factor is } v=e^{\int\limits p dx} \\ y'v+py v=q v \\ \text{ \choose } v'=vp \\ \text{ so we have } (yv)'=qv \\ \text{ then integrate both sides } \\ yv=\int\limits qv dx +C \\ y=\frac{1}{v} \int\limits qv dx +\frac{C}{v}\]

OpenStudy (freckles):

And earlier we said v'=vp so we have \[\frac{dv}{dx}=vp \\ \frac{dv}{v}=p dx \\ \int\limits \frac{dv}{v} = \int\limits p dx \\ \ln|v|=\int\limits pdx+C \text{ we will using this } C=0 \\ \ln|v|=\int\limits p dx \\ v=e^{\int\limits p dx}\]

OpenStudy (freckles):

like this is the method using the integrating factor

OpenStudy (anonymous):

thanks a lot!

OpenStudy (freckles):

where the solution is \[y=\frac{1}{v} \int\limits qv dx+\frac{C}{v} \\ \text{ where } v=e^{\int\limits pdx}\]

OpenStudy (anonymous):

you're brilliant! thank you!